5. You wonder whether the ratings for movies that are dramas differ, on average, from the ratings of movies that are comedies. You retrieve the ratings of 5 comedies and 5 dramas from the Internet Movie Database (IMDB). The ΣR for dramas is 25 and for comedies it is 31. What is the value of U for dramas?
= 25 + (25/2) – 25 = 12.5
= 25 + (30/2) – 25 = 15
= 25 + (30/2) – 31 = 9
= 25 + (25/2) – 31 = 18.5
Answer #1

Formula for U is as follows:

Image result for mann whitney u formula

Let 'a' represent Dramas and 'b' represent comedies

In this case na = nb = 5, Ra = 25 and Rb = 31

So, U for Dramas = Ua = 5*5 + 5*(5+1)/2 - 25

U for Dramas = Ua = 25 + 30/2 - 25 = 15 (2ndOption is correct)